OpenStudy (anonymous):

Choose the correct simplification of 3^8 over 3^3 243 15 1 over 243 1 over 15

OpenStudy (anonymous):

243... Since the base is 3 in both terms, just subtract the exponents. 3^8 / 3^3 = 3^5 = 243